OpenStudy (anonymous):

How do I simplify 2/x + 2/x - 1 - 2/x-2

OpenStudy (whpalmer4):

I assume that is \[\frac{2}{x}+\frac{2}{x-1}-\frac{2}{x-2}\]If you multiply each fraction by \[\frac{x(x-1)(x-2)}{x(x-1)(x-2)}\]you will get a common denominator after the dust settles.

OpenStudy (anonymous):

So I have to multiply each individual fraction by that?

OpenStudy (anonymous):

Multiply out each term by the denominator of the other terms. this will give you a single common denominator.